### v2

# MedS3: Towards Medical Small Language Models with Self-Evolved Slow Thinking

Abstract
Medical language models (MLMs) have become pivotal in advancing medical natural language processing. However, prior models that rely on pre-training or supervised fine-tuning often exhibit low data efficiency and limited practicality in real-world clinical applications. While OpenAI’s o1 highlights test-time scaling in mathematics, attempts to replicate this approach in medicine typically distill responses from GPT-series models to open-source models, focusing primarily on multiple-choice tasks. This strategy, though straightforward, neglects critical concerns like data privacy and realistic deployment in clinical settings. In this work, we present a deployable, small-scale medical reasoning system, MedS3, designed for long-chain reasoning in clinical tasks using a self-evolution paradigm. Starting with a seed dataset of around 8,000 instances spanning five domains and 16 datasets, we prompt a base policy model to perform Monte Carlo Tree Search (MCTS) to construct rule-verifiable reasoning chains. Each reasoning step is assigned an evolution rollout value, allowing verified trajectories to train the policy model and the process reward model (PRM). During inference, the policy model generates multiple responses, and the reward model selects the one with a newly proposed PRM-guided Vote-Sum (P-VS) strategy. Experiments on eleven evaluation datasets demonstrate that MedS3 outperforms not only the prior strongest medical model by 6.59, but also 32B-level general reasoning models by 8.71 points. Code and data are available at https://github.com/pixas/MedSSS.
MedS3: Towards Medical Small Language Models with Self-Evolved
Slow Thinking
Shuyang Jiang♣,♢, Yusheng Liao♠,♢, Zhe Chen♠,♢, Ya Zhang♠,♢, Yanfeng Wang♠,♢, Yu Wangthanks: Corresponding Author,♠,♢ Shanghai Jiao Tong University Fudan University Shanghai Artificial Intelligence Laboratory shuyangjiang23@m.fudan.edu.cn {liao20160907,chenzhe2018,ya_zhang,wangyanfeng622,yuwangsjtu}@sjtu.edu.cn
1 Introduction
Models Without Without Diverse Small Size Slow Thinking Process
Pretraining Close-sourced Teacher Clinical Coverage Reward Usage
MMed-Llama3
UltraMedical
HuatuoGPT-o1
O1-journey Part 3
MedS3
Table 1: Comparison of MedS3 with other medical models. Our MedS3 supports flexible inference-time scaling on resource-constrained devices, as well as process reward-guided decoding algorithms. Furthermore, MedS3 is a self-evolved model without dependence on large proprietary models for distillation or critique.
Large Language Models (LLMs) have demonstrated significant potential in the medical domain (Singhal et al., 2023; Nori et al., 2023; Chen et al., 2023b), with applications ranging from generating clinical notes (Biswas and Talukdar, 2024; Jung et al., 2024) to supporting patient communication (Tu et al., 2024; Liao et al., 2024b). Recently, slow-thinking reasoning models, exemplified by OpenAI o1 (OpenAI, 2024), have shown impressive improvements on reasoning-intensive mathematical problems (Lyu et al., 2025; Wang et al., 2024). However, limited efforts have been made to induce similarly strong reasoning abilities in medical-oriented language models.
Before the rise of OpenAI o1, researchers focused on improving medical language models (MLMs) primarily through extensive pre-training, which demands large computational resources and often yields only modest gains in downstream tasks. For instance, MMed-Llama3 post-pretrained Llama3 8B on 25.5B tokens but achieved only about a five-point performance gain compared to its base model (Qiu et al., 2024). Other works (Christophe et al., 2024; Ankit Pal, 2024) leverage supervised fine-tuning (SFT; Ouyang et al. (2022)) since it is computationally efficient. However, many human-annotated SFT datasets provide only concise responses (either short phrases or a single ground truth option), and hence fine-tuning on such data can degrade an MLM’s broader language fluency, reducing its practical value in realistic clinical scenarios. On the other hand, synthetic corpora generated by LLMs (Luo et al., 2024; Qiu et al., 2024) often contain hallucinations (Xu et al., 2024b; Huang et al., 2023), so directly using these outputs for teacher-forcing limits a target model’s optimization space. Consequently, focusing on inference-time scaling, often referred to as “slow thinking”, emerges as a data-efficient approach with the potential to enhance clinical mastery while mitigating the drawbacks of pre-training- or fine-tuning-heavy strategies.
A closely related work to slow thinking is HuatuoGPT-o1 (Chen et al., 2024), which employs GPT-4o to generate reasoning-intensive problems and corresponding complex reasoning steps for distillation and reinforcement learning (RL). While it achieves certain levels of long-chain reasoning, its heavy reliance on large proprietary models limits its generalizability to other clinical applications. Moreover, The adopted RL algorithm (Schulman et al., 2017) requires accurate value models, which is far from practicability in medicine yet.
Another relevant work, O1 Journey Part 3 (Huang et al., 2025), directly distills OpenAI o1’s outputs into $\sim$70B-parameter models. This approach compromises user-friendliness and data privacy, and also struggles to detect hallucinations. Moreover, relying exclusively on multiple-choice problems for distillation constrains its applicability to a broader range of clinical tasks.
In contrast, this work introduces MedS3, a small-scale medical language model equipped with robust long-chain reasoning capabilities (the “policy”) and an accompanying fine-grained process reward model (the “PRM”). Our approach uses a self-bootstrapping pipeline to enhance the model’s performance across diverse clinical tasks. Specifically, we curate 16 medical tasks from established medical training corpora, encompassing clinical diagnosis QA, medical natural language inference, knowledge-intensive QA, long-context QA, and biomedical QA. We then sample 500 seed instances from each task to form an initial dataset of approximately 8,000 instances, which kickstarts our self-evolution process. Leveraging a Monte-Carlo Tree Search (MCTS)-based evolution pipeline, we simultaneously generate synthetic datasets for both policy fine-tuning and PRM training. Rule-verified synthetic trajectories are preserved to optimize the policy model, where intermediate reasoning steps labeled with MCTS rollout values serve as the PRM dataset. By fine-tuning the base model on this enriched policy dataset and further enhancing it using our proposed PRM-guided Vote-Sum strategy, our system achieves holistically superior performance across eleven clinical reasoning benchmarks, surpassing both medical models and general reasoning models with much lower parameters.
The uniqueness of MedS3 is further underscored by a comprehensive comparison in Table 1, which demonstrates its superiority over other medical models in robust long-chain reasoning and breadth of clinical-task coverage. Overall, we summarize our contributions as follows:
  1. 1.
    First Self-Evolution Framework: To the best of our knowledge, this work introduces the first self-evolution framework specifically designed to empower small-scale medical models with long-chain reasoning capabilities, enabling data-efficient performance gains across a wide range of clinical applications.
  2. 2.
    State-of-the-Art Performance: Our self-evolved system, MedS3, achieves comprehensive performance improvements across eleven clinical benchmarks, outperforming all medical and large-sized reasoning models overall. This is driven by the integration of a fine-grained PRM that enhances reasoning accuracy at each step.
  3. 3.
    Open-Source Resources for Research: We openly release both the policy fine-tuning corpus and process reward model corpus, providing valuable resources for future research and fostering further advancements in medical AI.
2 MedS3
This section presents a detailed overview of the proposed MedS3 framework, which is structured into four key components:
  1. 1.
    Self-Bootstrapping Evolution (§2.1) which synthesizes reasoning trajectories as training data, with Monte-Carlo Tree Search (MCTS) technique using the base policy $\pi_{0}$.
  2. 2.
    Policy Model $\pi$ (§2.2) which is derived by fine-tuning the base policy $\pi_{0}$ using the generated synthetic data with supervised learning.
  3. 3.
    Process Reward Model (PRM) $V_{\theta}$ (§2.3) which is fine-tuned with step-wise supervision using soft-labels and assigns a value in the range $[0,1]$ to each reasoning step, representing correctness of individual steps.
  4. 4.
    PRM-guided Inference (§2.4) which utilizes the proposed PRM-guided Vote-Sum strategy to choose the final solution.
The overall framework is presented in Fig. 1.
Figure 1: Overview of the construction of MedS3 framework. MedS3 utilizes a Monte-Carlo Tree Search pipeline to self-generate step-by-step reasoning paths for each question in the seed dataset (a). During this process, MedS3 uses result simulation to obtain the rollout value for each node (b); After obtaining the child’s rollout value, MedS3 executes back-propagate to enable precise value prediction from deeper layer to transfer back to shallow nodes (c). After gathering all correct and wrong finish nodes, we use supervised fine-tuning to optimize the policy model $\pi$ with correct reasoning trajectories and step-wise discriminative loss to obtain a process reward model $V_{\theta}$ (d).
2.1 MCTS-guided Evolution
This algorithm builds upon an $n$-ary tree, where each tree node $T$ contains the following attributes: (1) reasoning step $s$, which is an intermediate step tracing from the root; (2) value $v$, which is an evaluation of the correctness of the current node; (3) the children nodes $\{c\}$, which is a collection of nodes that continue reasoning from the current node; (4) the parent node $p$ which is the former reasoning step and (5) the number of visits $n$. Every root node is initialized as $T=([s_{0}],0,\emptyset,\mathrm{null},0)$ where $s_{0}=$“Let s break down this problem step by step.” to guarantee a multi-step reasoning process. There are four stages in a full MCTS pipeline, including Node Selection, Node Expansion, Node Rollout, and Backpropagation.
Node Selection
This process starts from an initial root to select the next node. Within each iteration, we use UCB (Winands et al., 2008) as the criterion to select a child, which is as follows:
$UCB(T)=v_{C}+\gamma\sqrt{\frac{\ln n_{T_{parent}}}{n_{T}}}$ (1)
where $T_{parent}$ is the parent node of $T$ and $\gamma$ is an exploration constant set as 2. For each intermediate node, we select its child node with the highest $UCB$ value. We choose this criterion to expect models to further explore those nodes with high rollout values but low visiting counts, which inhibits models from repeatedly expanding already high-value nodes and encourages the expansion of underexplored nodes with the second highest value.
Node Expansion
After picking up the candidate node $T_{c}$ using the UCB criterion, we expand the reasoning steps of the current node. If the current node possesses a relatively high value ($v_{c}\geq thr$, where $thr=0.9$ is a pre-defined threshold), we prompt the node to directly generate a Finish node to accomplish this path reasoning. This manual operation not only does not impact the reasoning correctness as a value is close to 1 only when the trace $[s_{0},s_{1},\cdots,s_{k}]$ is close to the correct final answer, but also reduces unnecessary exploration tokens. Otherwise, assume that the selected node is located at $k$-th depth among the tree with previous reasoning trajectories $[s_{0},s_{1},\cdots,s_{k}]$ connected by a coherence phrase $t_{s}$, we sample $B$ single-step outputs $\{s_{k+1,i}\mid i=1,2,\cdots,B\}$ based on the previous trajectory using a Reason [^1] Prompts of Finish and Reason actions are illustrated in Appendix D node:
$s_{k+1,i}\sim\pi_{0}([s_{0}\oplus s_{1}\oplus\cdots\oplus s_{k}]\mid x)$ (2)
where $\oplus$ is the operation to connect two steps using the coherence phrase $t_{s}$, $\pi_{0}$ is the base policy model, $x$ is the original input prompt and $s_{k+1,i}$ is sampled with a high temperature (1.0) to harvest a diverse search space. To control that each reasoning step is atomic enough for the whole thought path, we set the stop tokens as Step k+2. Subsequently, the $B$ nodes $\{T_{i}=([s_{0},s_{1},\cdots,s_{k},s_{k+1,i}],0,\emptyset,T_{c},0)\mid i=1,2,\cdots,B\}$ are added to $T_{c}$ as children nodes.
Node Rollout
As the PRM is not yet available during evolution, the rollout process is conducted using simulation to obtain the estimated value for a chosen node. Specifically, for a chosen unvisited node $T_{c}$, we set a simulation budget $L=\min(L_{\min},\frac{L_{0}}{k})$ where $k$ is reasoning step counts, to encourage sufficient simulation trials when the known reasoning path is short, but expect to see a deterministic reasoning result conditioning on a long trajectory. After that, we prompt the policy model $\pi_{0}$ to directly output the answer $L$ times:
$a_{c}^{l}\sim\pi(x,[s_{0},s_{1},\cdots,s_{k}])$ (3)
where $l\in[1,L]$ and $a_{c}^{l}$ is the $l$-th simulated answer. The accuracy of the $L$ simulations $acc=\frac{1}{L}\sum_{l=1}^{L}\mathds{1}_{a_{c}^{l}=y}$ is assigned as the value of $T_{c}$. For the Finish node, the rollout value is computed by comparing the derived answer and the ground truth and an $0/1$ value is assigned.
Backpropagation
After obtaining the value for the selected node, we conduct value back-propagation starting from $T_{c}$ till the root node, updating all tree node values among the trace. This process aims to use more accurate value estimation in the deeper layer to update early nodes’ values so that these values serving as PRM’s optimization labels become synchronously precise. Specifically, for an arbitrary node $T_{k}$, we propose to update its visits $n_{k}$ and $v_{k}$ as follows:
$n_{k}$ $=n_{k}+1$ (4)
$v_{k}$ $=\frac{1}{2}\left(v_{k}+\frac{\sum_{ch}v_{ch}\cdot n_{ch}}{\sum_{ch}n_{ch}}\right)$ (5)
Note that this update process considers both its children values and its own value, which guarantees that the parent node will never have the same value as its children. It considers both correctness and completeness for the evaluation of a reasoning step.
Termination of Search
Due to computational limits, we cannot afford excessive computing resources to fully expand the tree for a training instance. Therefore, we set two criteria to terminate the search process. Firstly, we set the minimum correct nodes to $\tau$. Once the total correct count in the tree exceeds $\tau$, we stop the exploration of this tree. Second, if there are no correct nodes after affording a certain number of node exploration trials, we prompt $\pi_{0}$ to generate Finish node for all leaves. The first strategy aims to reduce extra computation, while the second expects to obtain as more correct answers as possible, for the optimization of both policy and reward models.
2.2 Policy Model Fine-tuning
The policy training mainly leverages the correct nodes $T_{k}^{1}$ which are assigned a value $1.0$ during rollout and corresponding reasoning trajectories gathered before: $D_{\pi}=\{(T_{k}^{1},[s_{0}\oplus s_{1}\oplus\cdots\oplus s_{k}])\}$. These correct reasoning traces are supervised fine-tuned to deduce a self-improved policy model:
$\mathcal{L}_{\pi}=\frac{1}{L_{k}}\sum_{i=1}^{L_{k}}-\log p_{\pi}(y_{i}|x,y_{<i})$ (6)
where $y_{i}$ is the $i$-th token of the reason trajectory and $L_{k}$ is the total length of the trajectory.
2.3 PRM Fine-tuning
Dataset Collection
We gather all finish nodes and their corresponding traces: $\{(T_{k}^{j},[s_{0},s_{1},\cdots,s_{k}^{j}],v_{k}^{j})\mid k=1,2,\cdots,K;j=1,2\cdots,J\}$, where $K$ is the max depth of the evolved tree, $J$ is the max width of the $j$-th layer, and $v_{k}^{j}\in(0,1)$. The reasoning traces $[s_{0},s_{1},\cdots,s_{k}^{j}]$ are assembled with $t_{s}$. This naturally enables $V_{\theta}$ to distinguish between siblings inheriting from the same parent node: $T_{k}^{j_{1}}$ and $T_{k}^{j_{2}}$ with distinct rollout values $v_{k}^{j_{1}}$ and $v_{k}^{j_{2}}$. To prevent $V_{\theta}$ from being biased by the distribution of value 0 and value 1, we tallied the correct and incorrect finish nodes: $\{T_{k}^{0}\mid v_{T_{k}}=0\}$ and $\{T_{k}^{1}\mid v_{T_{k}}=1\}$, and randomly sampled elements $\min\left(|\{T_{k}^{0}\}|,|\{T_{k}^{1}\}|\right)$ from these two sets, maintaining a balanced distribution of the nodes of correct and incorrect reasoning: $\{\hat{T}_{k}^{0}\}$ and $\{\hat{T}_{k}^{1}\}$. The sampled nodes are combined and form the final $V_{\theta}$ tuning set: $D_{V_{\theta}}=\{\hat{T}_{k}^{0}\}\cup\{\hat{T}_{k}^{1}\}$.
Learning objective
Previous works in the math domain choose to directly learn the rollout value (Zhang et al., 2024a) or learn the pair-wise ranking preference (Guan et al., 2025). However, in our work, we choose to learn the prediction of the correctness probability of an intermediate step using a 2-class cross-entropy loss. The PRM $V_{\theta}$ is tuned based on the tuned policy model, with the language model head replaced by a token classification layer with a cross-entropy loss at the end of each step. Although Zhang et al. (2025) suggests that the PRM label should be set to 1 (a hard label) once the rollout score is above zero, empirically we borrow the insight from label-smoothing (Szegedy et al., 2016), where a soft label groups positive internal steps tightly in the representation space. We show the comparison between the soft and hard label in Appendix B.1. Specifically, we set the label $y_{k}$ as the original node rollout value, and optimize $V_{\theta}$ using the following loss function:
$\mathcal{L}_{V_{\theta}}=\frac{1}{|D_{V_{\theta}}|}\sum_{T_{k}\in D_{V_{\theta}}}y_{k}\log\hat{y}_{k}+(1-y_{k})\log(1-\hat{y}_{k})$ (7)
where $\hat{y}_{k}$ is the predicted probability of the given step and $y_{k}$ is the label. This soft-label training, not only encourages $V_{\theta}$ to cluster preferred and dis-preferred steps, but also prevents the learning of fuzzy labels (rollout value around 0.5).
Figure 2: Overview of the PRM-guided Vote-Sum.
2.4 PRM-guided Decoding
Referring to the inference scaling law (Wu et al., 2024d), enlarging the inference token budget is a deterministic way to enhance the downstream reasoning performance. To balance between the advantages of enlarging the inference budget and unwillingly reduced efficiency, we propose a new decoding strategy named PRM guided Vote-Sum that considers both response estimation and semantic coherence, which is illustrated in Fig. 2.
PRM guided Vote-Sum (P-VS)
This method employs $\pi$ to sample $M$ candidate responses $\{y_{m}\mid m=1,2,\cdots,M\}$ given an input problem $x$ and uses the PRM $V_{\theta}$ to select the response whose answer is estimated to have the highest values in total. Specifically, for a simple output $y_{m}$, we split it into $K$ steps $\{s_{m}^{k}\mid k=1,2,\cdots,K\}$ with pre-defined coherence phrase $t_{s}$. The PRM assigns a score $q\in[0,1]$ for each internal step $\{q_{m}^{k}=V_{\theta}(s_{m}^{1}\oplus s_{m}^{2}\oplus\cdots\oplus s_{m}^{k})\mid k=1,2,\cdots,K\}$. The overall response score $v_{y_{m}}$ takes the minimum value of the score chain (Lightman et al., 2023) or the last value (Zhang et al., 2025). The Vote-Sum strategy comprehensively considers the occurrence of semantically equivalent outputs but also the confidence score predicted by $V_{\theta}$:
$a_{v}$ =$=\{(a_{1},v_{a_{1}}),(a_{2},v_{a_{2}}),\cdots,(a_{n},v_{a_{n}})\}$ (8)
$y_{target}$ $=\operatorname*{argmax}_{y_{m}\in\{y_{j}\mid a_{y_{j}}=\operatorname*{argmax}_{a_{i}}v_{a_{i}},(a_{i},v_{a_{i}})\in a_{v}\}}v_{y_{m}}$ (9)
where $n\leq M$ is the number of different answers obtained from the $M$ responses, $v_{a_{n}}=\sum_{m}\mathds{1}_{a_{y_{m}}=a_{n}}\cdot v_{y_{m}}$ is the sum of values of samples whose answers equal to $a_{n}$.
Figure 3: Overview of the used seed datasets.
3 Data Statistics
A slow-thinking system in medical scenarios should both excel at exam-level question answering (QA) and handling real-world clinical scenarios, like diagnosis (Tchango et al., 2022), specific disease syndrome (Lab, 2020) and drug-related problems (Huynh et al., 2016). However, previous works only focused on a simple scenario, with only limited data diversity, especially multiple-choice QA, to train reasoning models. To approximate realistic clinical usage and promote medical reasoning models on a diverse range of clinical tasks, we curate a training corpus, from 16 existing public medical datasets and divide them into five dimensions according to the task category. For each dataset, we randomly select 500 items and form a seed dataset with around 8,000 instances. We show the dataset statistics in Fig. 3. The details about the definition of the five dimensions and the corresponding tasks can be found in Appendix E.
Models Knowledge QA BioMedical QA Long Context QA NLI Diagnosis QA Avg.
MedQA MedMCQA PubMedQA BioASQ MMLU BioMRC DDXPlus PubHealth HealthFact DrugDose SEER
Proprietary language models
GPT-4o-mini 75.81 67.58 47.80 83.01 83.79 66.85 54.00 59.14 65.24 73.91 54.54 66.52
GPT-3.5-turbo 59.31 58.12 37.40 74.11 71.11 56.22 39.05 57.84 67.85 86.96 73.61 61.96
Open-source language models
Qwen2.5-7B 55.54 54.12 53.40 73.62 74.38 56.48 31.25 57.11 52.69 60.87 33.07 54.78
Llama3-8B 57.50 55.92 56.40 75.73 68.55 56.50 35.30 64.09 70.88 73.91 47.07 60.17
Llama3.1-8B 61.51 57.42 59.00 71.36 72.52 55.60 19.00 61.82 63.97 73.91 52.62 58.98
R1-Distill-Llama8B 50.12 48.89 46.60 70.55 68.42 53.49 36.10 55.73 62.04 69.57 31.71 53.93
QwQ-32B-preview 68.89 61.03 48.60 73.62 74.18 79.76 45.40 63.36 66.08 39.13 37.26 59.76
R1-Distill-Qwen32B 76.83 66.27 38.20 78.32 85.07 78.66 53.90 59.95 63.80 82.61 26.22 64.53
Open-source medical models
MMedS-Ins 53.57 48.24 56.60 77.35 50.86 31.47 97.53 54.26 69.64 95.65 97.93 66.65
MedLlama3 55.85 59.36 66.40 84.63 70.08 47.97 22.50 62.39 68.10 69.57 50.69 59.78
Med42 50.20 49.70 55.40 74.76 61.43 57.26 31.35 59.14 81.57 65.22 37.14 56.65
OpenBioLLM 50.20 50.56 41.40 47.73 61.69 27.46 16.55 18.77 53.28 34.78 46.48 40.81
UltraMedical3-8B 68.89 61.82 51.60 80.58 75.08 45.18 36.70 66.13 72.73 60.87 24.55 58.56
UltraMedical3.1-8B 70.93 62.78 56.40 77.18 76.43 54.26 31.55 59.14 70.20 56.52 45.86 60.11
Open-source slow-thinking medical models
HuatuoGPT-o1 62.53 59.31 69.20 87.70 70.53 50.98 40.20 24.61 66.08 56.52 46.85 57.68
MedS3 (Ours)
    CoT 65.91 60.55 56.80 78.48 75.66 55.84 51.65 57.03 64.73 73.91 48.97 62.68
    SC 70.93 64.21 58.20 79.13 79.63 63.66 57.00 64.42 70.37 86.96 52.19 67.88
    P-VS 71.88 65.20 59.60 80.10 79.50 77.12 65.20 73.03 79.97 95.65 58.36 73.24
Table 2: Experiment results in 11 medical datasets among four types of models. We highlight the best results with bold and underlines the second-best results. MedS3 with PRM guided Vote Sum (P-VS) achieves superior performances on real-world clinical datasets.
4 Experiments
In this section, we comprehensively evaluate the proposed MedS3 on various downstream domains, including both in-domain and out-of-domain datasets.
4.1 Experiment Setups
Training and Evaluation
We choose Llama3.1-8B-Instruct as the initialization of both the policy model and PRM model. We select MedQA (Jin et al., 2021), PubMedQA (Jin et al., 2019), MedMCQA (Pal et al., 2022), PubHealth (Kotonya and Toni, 2020), BioMRC (Pappas et al., 2020), HealFact Classification (Kotonya and Toni, 2020), Drug Dose Extraction (Huynh et al., 2016), DDX-Plus (Tchango et al., 2022) as the in-domain evaluation benchmarks, the medical parts of MMLU (Hendrycks et al., 2021), BioASQ (Tsatsaronis et al., 2012) and SEER Classification (Dubey et al., 2023) as the out-of-domain evaluation sets. We provide three different decoding strategies for MedS3, including CoT (Wei et al., 2022), Self-Consistency (SC) and our proposed P-VS. The hyperparameters of synthesis, self-training and evaluation are presented in Appendix F.
Baselines
We choose the following categories to serve as baselines: (1) Proprietary general LLMs, including GPT-3.5-turbo (OpenAI, 2022) and GPT-4o-mini (OpenAI, 2023); (2) Open-sourced general LLMs, including Llama 3 8B, Llama 3.1 8B (Dubey et al., 2024) and Qwen2.5 7B (Yang et al., 2024), QWQ-preview-32B (Team, 2024) and two distilled models from DeepSeek R1 (Guo et al., 2025): R1-Distill-Llama8B and R1-Distill-Qwen32B; (3) Open-sourced Medical LLMs, including MedLlama 3 8B [^2] https://huggingface.co/ProbeMedicalYonseiMAILab/medllama3-v20, MMedS-Ins-Llama-3-8B (Wu et al., 2024b), Med42 (Christophe et al., 2024), OpenBioLLM (Ankit Pal, 2024), and UltraMedical3-8B and UltraMedical3.1-8B (Zhang et al., 2024b). We also compare our method with HuatuoGPT-o1-8B (Chen et al., 2024).
4.2 Main Results
We present the comprehensive experiment results in Table 2. The results unveil that most prior medical LLMs show superior results in traditional medical benchmarks (MedQA or PubMedQA); while such superiority cannot generalize to out-of-distribution real-world clinical benchmarks (DDXPlus or DrugDose), which results in their sub-optimal overall performance compared to Llama3-8B. In contrast, our MedS3 is not optimized exclusively for multiple-choice medical datasets and hence achieves the best performance among all open-sourced competitions. As an 8B system, MedS3 achieves +13.07 average performance gains with respect to the base model in the overall assessment, which not only outperforms medical-oriented models as well as general reasoning models.
Specifically, compared to HuatuoGPT-o1 and MedLlama3, MedS3 shows superior performance on reasoning-intensive benchmarks, including MedQA and MedMCQA, as well as clinical benchmarks. This verifies that MedS3 learns medical reasoning philosophies and the clinical deduction process. Another model MMedS-Ins, which directly post-pretrains and fine-tunes on millions of clinical corpus, harvests superior performance on in-domain clinical tasks, such as SEER and DDXPlus. However, directly fine-tuning on question-answer pairs inevitably makes the model lose the ability to output long responses, which is extremely important in reasoning tasks. In contrast, our model MedS3, possesses a comprehensive performance on both traditional tasks and clinical scenarios with strong reasoning abilities.
5 Analysis
Select Model Value Estimation Traditional QA Clinical Tasks
Majority ORM PRM Min Prod Last Vote-Sum MedQA MedMCQA Healthfact DDXPlus SEER Mean
70.93 64.21 70.37 57.00 52.19 62.94
67.87 60.63 63.38 53.75 48.71 58.87
71.88 64.67 70.12 57.15 53.44 63.45
67.64 59.20 73.57 52.40 41.24 58.81
65.51 60.00 75.51 57.30 47.67 61.20
60.17 58.00 78.37 65.10 61.67 64.66
71.88 65.20 79.97 65.20 58.36 68.12
Table 3: Comparison of P-VS with other decoding methods under the same token budgets. “Majority” means we use SC to select the final response. “ORM with Last” means selecting the response with the max ORM value.
5.1 Effectiveness of P-VS
In this section, we compare our proposed decoding method P-VS to the previously widely adopted PRM decoding method: Best-of-N (BoN) as well as ORM guided methods (ORM with Last and ORM with Vote-Sum). We compare three variants of BoN, including BoN-min (Lightman et al., 2023) which takes the minimum step value of a reasoning trajectory as the estimation of the whole sequence, BoN-prod (Lightman et al., 2023) which takes the production of step values instead, and BoN-last (Zhang et al., 2025), which treats the PRM as an ORM. We select two traditional QA tasks and three clinical tasks to conduct comparison and show results in Table 3. A very interesting finding is that in the two traditional QA datasets, BoN-last achieves extremely bad results, but shows great performance on par with Vote-Sum in clinical tasks. We hypothesize that clinical tasks and Olympia-level math problems are both difficult for Llama3.1-8B, and hence the findings that PRM behaves more like an ORM transfer successfully to clinical tasks. On the other hand, on relatively advantageous tasks, MC-rollout value encodes values of both future and current steps, which leads the PRM to perform as a human-labeled PRM to distinguish steps precisely. As a result, in traditional medical tasks, we use the minimum value among step values to represent the whole sequence’s correctness, while in clinical tasks, we use the last step’s value instead. Compared to ORM-based methods, they only excel at traditional QA benchmarks, lagging far behind all PRM-based methods in clinical usage, further validating the efficacy of process supervision. ORM with Vote-Sum surpasses ORM with last by a large margin, which also verifies the effectiveness of the proposed Vote-Sum strategy. We present the error rate of PRM in Appendix B.2 to show that rule-verified PRM is sufficient in detecting erroneous clinical reasoning steps.
Figure 4: Budget scaling on clinical benchmarks. When the token budget increases, MedS3’s improves rapidly and then slows down as the budget is further increased.
5.2 Inference-Time Scaling
In this section, we discuss whether MedS3 can benefit from more inference tokens. We sample $n=2,4,8,12,16$ candidates for a given problem, and conduct P-VS to select the most plausible response. We select five clinical datasets (Healthfact, DDXPlus, SEER, BioMRC, Pubhealth) to illustrate the scaling law, and plot the results in Fig. 4. We observe that the first 4 samples provide significant performance improvements, showing a satisfactory tradeoff between the tokens consumption and the performance improvement. Although more generations bring fewer significant performance gains, the increasing trend never slows down except for the Pubhealth dataset, illustrating nearly unbounded scaling potentials.
Figure 5: Three widely adopted methods to empower models with medical reasoning abilities. MCTS+PRM is the best among the three, making it the core of MedS3.
5.3 Comparison of Slow-Thinking Styles
In this section, we compare MCTS plus PRM which is what MedS3 leverages, with distillation from strong reasoning models, which is what O1-journey-part3 (Huang et al., 2025) does and pure reinforcement learning (RL), which is what DeepSeek-R1 (Guo et al., 2025) adopts. These methodologies are widely used for empowering small language models with strong long-chain reasoning abilities. We use the same dataset in §3 to implement RL, and use the officially released distillation dataset provided by Huang et al. (2025) to SFT the base model. The results presented in Fig. 5 demonstrate that in exam-level medical QA datasets where the base model already excels at, distillation from large proprietary reasoning models is much more data-efficient than the other two methods, albeit sacrificing generalization in clinical tasks. In contrast, with both considerable performance leap and generalization, RL is second to MCTS+PRM. We hypothesize that the medical diagnosis step is easier to determine than math reasoning steps, resulting in a more accurate PRM.
6 Conclusion
In this paper, we present MedS3, a self-evolved slow-thinking medical language model built for universal clinical usage. We collect a seed dataset, with only 7,465 instances but covering 16 different realistic medical tasks, and use Monte-Carlo Tree Search to construct policy data and PRM data. We propose a new decoding method, which enables the resulting policy model to collaborate with the fine-tuned PRM model, to produce credible long-chain responses. Experiment results demonstrate that our model achieves superior performance on eleven downstream medical benchmarks, especially in realistic clinical ones, surpassing open-sourced models with a large margin.
Limitations
MedS3 achieves superior performance over eleven benchmarks by conducting MCTS in seed datasets to collect both policy and PRM training data and a newly proposed decoding strategy: PRM-guided Vote-Sum. However, it can be further improved via these strategies: (1) conduct an iterative self-evolution pipeline to enhance both the policy and PRM; (2) cooperate with reinforcement learning to empower the policy with the “aha-moment” (Guo et al., 2025) ability; (3) introduce more training samples to avoid data imbalance.
Ethics Considerations
In developing clinical slow-thinking model MedS3, it is crucial to address ethical consideration that arise when utilizing AI in healthcare environments. Below are the key ethical considerations that have been taken into account:
Perfomance vs. Potential Risks
While MedS3 demonstrates significant enhancements in clinical reasoning and task performance, it is important to acknowledge the inherent limitations of AI models. These models can generate misleading information of “hallucinations”, which could pose risks in clinical settings. Therefore, MedS3 is not intended to replace medical professionals or provide definitive clinical decisions but rather to assist healthcare providers under appropriate supervision.
Data Ethics and Privacy Compliance
All patient data has been anonymized, and informed consent was obtained for its use, ensuring full compliance with privacy policies and obtaining explicit permission for all data usage. Additionally, data usage has been approved by relevant ethics committees, ensuring compliance with ethical standards and privacy protection requirements.
Appendix A Related Works
With the success of the generalist LLMs, their advancements in both clinical and biomedical scenarios have shown significant promise. Numerous previous works focus on developing medical-specific LLMs, which are now being increasingly adopted across various clinical settings. These medical LLMs generally follow three main approaches:
Continual Pre-training Medical LLMs
These types of medical LLMs Xu (2023); Chen et al. (2023c) are developed on the advantage of generalist LLM and attempt to inject domain-specific knowledge and expertise through continual pre-training techniques. Such type of methods usually require significant computational resources. For example, 3 billion training tokens are used in HuatuoGPT-II (Chen et al., 2023a) and PMC-Llama (Wu et al., 2024a) even requires more than 75 billion tokens. However, results in recent works (Qiu et al., 2024) show that the benefits of continued pre-training are diminishing as the capabilities of the generalist LLMs improve.
Fine-tuned Medical LLMs
Compared to continuous pre-training, fine-tuning is a more efficient approach. It can rapidly adapt to medical scenarios and perform the relevant tasks effectively when the base LLMs are sufficiently powerful. (Ankit Pal, 2024; Christophe et al., 2024; Zhang et al., 2024b) Specifically, Liao et al. (2024a) develops a two-stage method which can decouple the knowledge-injection and clinical alignment procedure during the fine-tuning process to prevent the ‘alignment-tax.’ Wu et al. (2024c) collects a wide range of medical language processing tasks spanning 19 task categories and 122 unique clinical tasks to improve the LLMs’ capacities on various downstream clinical tasks.
Slow-Thinking Medical LLMs
With the significant achievements of the o1 (Jaech et al., 2024) in complex reasoning tasks, previous works show the potential advantage of the o1-like models in medical tasks (Xie et al., 2024; Xu et al., 2024a; Nori et al., 2024). Based on these, previous works develop the slow-thinking medical LLMs with distillation: Huang et al. (2025) directly learn the reasoning trajectory generated by o1 and Chen et al. (2024) improving the model’s reasoning ability through o1 synthesis of reflective data and reinforcement learning. Besides, Yu et al. (2025) create a Chinese version slow-thinking medical LLMs by constructing the preference data with QwQ (Team, 2024).
Appendix B Further Experiments
In this section, we present more experiments to validate the effectiveness of MedS3.
B.1 Soft Train vs. Hard Train
We here compare the PRM training techniques, where the step label is set to a hard two-class $0/1$ label (Zhang et al., 2025) or a soft float label introduced in MedS3. We remain the other settings unchanged and set the label to $1$ if the step’s rollout value is greater than 0, otherwise 0 when training the hard PRM. The results in Table 4 indicate that soft-label is comprehensively superior to hard-label. Although Zhang et al. (2025) indicate that hard-label works after prompting an LLM to filter bad intermediate steps, for the sake of privacy in clinical uasge and fairness we do not conduct filtering.
B.2 Reliability of PRM
Although our process reward model (PRM) is trained using rollout-value—a rule-based supervision signal—we empirically demonstrate that the PRM exhibits strong capability in identifying erroneous reasoning steps. To evaluate this, we compare two metrics: the task error rate (defined as instances where model predictions deviate from ground truth answers) and the PRM error rate (occurring when the PRM assigns a higher score to an incorrect candidate than to the ground truth-aligned prediction). As illustrated in Table 5, the majority of tasks exhibit a PRM error rate of zero, with the exception of the Healthfact dataset. This suggests that our fine-tuned PRM effectively differentiates between valid and invalid clinical reasoning steps. Furthermore, the observed discrepancy between task and PRM error rates implies that most errors arise not from the PRM’s assessment but from the policy model’s failure to generate candidates aligned with the ground truth. This finding underscores the need to refine the policy model in future iterations to address this limitation.
Method MedQA MedMCQA Healthfact DDXPlus SEER Average
Hard 64.81 62.18 77.61 64.35 54.94 64.778
Soft 71.88 65.20 79.97 65.20 58.36 68.122
Table 4: Comparison between soft-label which is adopted in MedS3 and hard-label.
Error Type
MedQA
MedMCQA HealthFact PubmedQA BioASQ Med MMLU BioMRC DDXPlus PubHealth DrugDose SEER
Task 28.91 34.60 20.03 40.40 23.79 20.50 22.88 34.60 26.97 4.35 41.87
PRM 0.00 0.00 16.16 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Table 5: Comparison of error rates (lower is better) in task-level and PRM-level, where PRM works with no errors in most of testbeds.
Appendix C Future Work
As a pioneering work, we have validated that small language models can self-evolve to empower themselves with strong reasoning abilities in clinical usage. There are several remaining directions to further enhance MedS3:
  1. 1.
    Conduct iterative evolution. Currently we are only focusing on one iteration evolution, which greatly leaves the PRM under-tuned.
  2. 2.
    Conduct Human-interference evaluation. MC-rollout value is verified to be not the best choice for evaluating the value of an internal step. We are eager to introduce fine-grained step label to enhance the optimization of the PRM.
  3. 3.
    Introduce more clinical data, not limited to close-ended generation. Currently, all the data used in MedS3 are close-ended, and the application of reasoning is not limited to such a narrow room. We intend to extend MedS3 to broader clinical tasks to make MedS3 a more useful system.
We will continue our exploration and make MedS3 more practical in medical domains.
Appendix D Prompt Template
We show the prompt used to synthesize reasoning data in Fig. 6 and Fig. 7.
Reason Template <|begin_of_text|><|start_header_id|>system<|end_header_id|>
Cutting Knowledge Date: December 2023 Today Date: 23 July 2024
<|eot_id|><|start_header_id|>user<|end_header_id|>
Reasoning Example: {Few-shot Example}
You are a professional medical expert majored at reasoning in hard medical-related problems.
Think critically about the problem and answer with concise, accurate reasoning. Please ensure your reasoning is thorough and elaborate, breaking down each step of your thought process.
Problem: {problem}<|eot_id|><|start_header_id|>assistant<|end_header_id|>
Step 0: Let’s break down this problem step by step
Step 1:
Figure 6: Reason template
Finish Template <|begin_of_text|><|start_header_id|>system<|end_header_id|>
Cutting Knowledge Date: December 2023 Today Date: 23 July 2024
<|eot_id|><|start_header_id|>user<|end_header_id|>
Reasoning Example: {Few-shot Example}
You are a professional medical expert majored at reasoning in hard medical-related problems.
Use thorough and elaborate steps to complete your reasoning. Conclude the task by stating: "The answer is {answer}".
Problem: {problem}<|eot_id|><|start_header_id|>assistant<|end_header_id|>
Step 0: Let’s break down this problem step by step
Step 1:
Figure 7: Finish template
Appendix E Dataset Details
In this section, we elucidate the seed dataset and the evaluation sets. We divide the used 16 training datasets into the following five dimensions:
  1. 1.
    Long Context QA: This dimension enables MedS3 to capture useful information from the given context and response with long-chain reasoning. This dimension covers BioMRC (Pappas et al., 2020), HeadQA Topic Classification (Vilares and Gómez-Rodríguez, 2019; Wu et al., 2024b), and DDX-Plus (Tchango et al., 2022).
  2. 2.
    Knowledge-Intensive QA: This dimension teaches MedS3 to use long-chain reasoning to answer knowledge-intensive problems, which covers MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), PubMedQA (Jin et al., 2019) and PubHealth (Kotonya and Toni, 2020).
  3. 3.
    Bio-Medical QA: This part leverages general data in bio-medicine domains to enhance the generality of MedS3, which includes SciQ (Welbl et al., 2017), Evidence Inference (DeYoung et al., 2020) and Head QA (Vilares and Gómez-Rodríguez, 2019).
  4. 4.
    Medical Natural Language Inference: This dimension prompts MedS3 to discriminate biomedical research concepts and corresponding descriptions, which contains Healthfact Classification (Kotonya and Toni, 2020), Medical Question Pair (MQP; McCreery et al. (2020)), and catalonia-independence-corpus (CIC; Zotova et al. (2020)).
  5. 5.
    Diagnosis QA: This dimension is related to real-world clinical scenarios, including disease diagnosis and classification and drug related questions. We choose Covid-19 Classification (Lab, 2020), Drug-Dose Extraction and Adverse Drug Event Classification (Huynh et al., 2016; Wu et al., 2024b).
The descriptions of each training and evaluation datasets are presented below:
  1. 1.
    MedQA (Jin et al., 2021) is a widely used benchmark for evaluating AI systems in medical question answering, featuring multiple-choice questions from professional medical licensing exams such as the USMLE and exams from China and Taiwan. We adopt its 5-options English version, taking its training set as seed data and 1,273 test problems as the evaluation benchmark.
  2. 2.
    PubmedQA (Jin et al., 2019) is a specialized benchmark for biomedical question answering, consisting of question-answer pairs derived from PubMed abstracts. It focuses on yes/no/maybe questions that require reasoning over biomedical literature. We use the human-labeled question set and split the training set and test set, with both 500 problems for evolution and evaluation, respectively. Note that we do not include relevant contexts before questions, challenging models’ internal knowledge comprehension.
  3. 3.
    MedMCQA (Pal et al., 2022) is a large-scale benchmark for medical question answering, featuring over 194,000 multiple-choice questions sourced from Indian medical entrance exams and other educational resources. It spans a wide range of medical topics, including anatomy, pharmacology, and pathology, and is designed to evaluate the reasoning and knowledge application skills of AI systems in a clinical context. The test set contains 4,183 problems.
  4. 4.
    MMLU (Hendrycks et al., 2021) is to measure LLM’s multitask accuracy, which contains 14,421 problems. The test covers 57 tasks including elementary mathematics, US history, computer science, law, and more. We select its medical-related problems, resulting in a test set with 1,561 problems.
  5. 5.
    BioMRC (Pappas et al., 2020) is a collection of medical-related question-answer pairs, specifically designed for the evaluation of machine reading comprehension (MRC) tasks in the biomedical domain. It is derived from a wide range of medical texts, including clinical notes, research papers, and medical textbooks. The dataset contains a series of questions and corresponding answers, where the answers are extracted from relevant passages of text. We use its 6,250 test set as the evaluation set.
  6. 6.
    HeadQA (Vilares and Gómez-Rodríguez, 2019) is a specialized medical question-answering dataset designed to evaluate models in the context of neurology and head-related disorders. It consists of a collection of questions paired with answers derived from a variety of clinical notes, medical reports, and other head-related health data sources.
  7. 7.
    DDX-Plus (Tchango et al., 2022) is a comprehensive medical diagnostic dataset designed to assist in the development and evaluation of machine learning models for differential diagnosis in clinical settings. It consists of clinical cases, where each case includes a set of symptoms, patient history, physical examination findings, and diagnostic questions, along with a list of potential diagnoses ranked by their likelihood. The diverse set of cases in the dataset spans multiple medical specialties, making it an ideal resource for creating models capable of assisting healthcare professionals in making informed diagnostic decisions. Due to its huge test set (over 100,000 test instances), we randomly select 2,000 items for evaluation.
  8. 8.
    SciQ (Welbl et al., 2017) is a scientific question-answering dataset designed to assess machine learning models in answering factual questions across a wide range of scientific domains. It consists of over 13,000 questions derived from scientific literature, including topics in physics, biology, chemistry, and earth sciences, among others. Each question is paired with a correct answer and is supported by a passage of text from which the answer is extracted.
  9. 9.
    Evidence Inference (DeYoung et al., 2020) is a collection designed to evaluate machine learning models on their ability to infer logical conclusions from evidence presented in the form of textual information. This dataset consists of structured pairs of premises (evidence) and hypotheses, where the goal is for models to determine the logical relationship between them—whether the hypothesis is supported, contradicted, or is neutral with respect to the provided evidence. Typically used for tasks such as textual entailment or natural language inference (NLI), the dataset includes a variety of complex scenarios across multiple domains, including law, healthcare, and science, where reasoning based on available evidence is crucial.
  10. 10.
    PubHealth (Kotonya and Toni, 2020) is a comprehensive dataset for explainable automated fact-checking of public health claims. Each instance in the PUBHEALTH dataset has an associated veracity label (true, false, unproven, mixture). Furthermore each instance in the dataset has an explanation text field. The explanation is a justification for which the claim has been assigned a particular veracity label. We construct two different test sets. Healthfact is to directly predict whether a given instance is true/false/unproven/mixture. The other, Pubhealth, is to predict whether the instance sentence and the given explanation expresses the same meaning.
  11. 11.
    Medical Question Pair (McCreery et al., 2020) contains a dataset of 3,048 similar and dissimilar medical question pairs hand-generated and labeled by Curai’s doctors. Models should clarify whether the given two questions are similar or not.
  12. 12.
    Catalonia-independence-Corpus (Zotova et al., 2020) is a dataset built for stance detection in Twitter for the Catalan and Spanish languages, with the aim of facilitating research on stance detection in multilingual and cross-lingual settings.
  13. 13.
    Covid-19 Classification (Lab, 2020) is an extension of the Hedwig library and contains all necessary code to reproduce the results of some document classification models on a COVID-19 dataset created from the LitCovid collection.
  14. 14.
    Adverse Drug Event (Huynh et al., 2016) is critical for developing automated systems that can support clinicians in identifying harmful drug reactions, potentially reducing healthcare costs, and enhancing patient safety. Given the increasing volume of clinical data, this dataset plays a key role in advancing AI-driven drug safety research and improving the overall quality of healthcare. We build Drugdose extraction test set to benchmark models to extract the exact dose of a specific drug.
Appendix F Hyperparameters
F.1 Data Synthesis
[…]
F.2 Self-Training of Policy and PRM
Policy tuning
We use trl [^3] https://huggingface.co/docs/trl/index as the training framework. We use vanilla SFTTrainer to train the policy model. We set the warmup ratio to 0.03 and the max sequence length to 8192. The batch size is set to 64 and the learning rate is set to 2e-4. We use LoRA (Hu et al., 2022) to efficiently train the model, where the rank $r$ and alpha $\alpha$ are set to 16 and 32, respectively. We train 1 epoch for the total 24,441 positive synthesized samples.
PRM tuning
We use PRMTrainer of trl to train the PRM model. We use the same hyper-parameters as those used in the policy fine-tuning but conduct a simple data filtering method. For an instance where all its solution trajectories are all correct or incorrect, we only randomly maintain one of them, as the excessively simple or hard instance will disturb the training process. For other instances, we guarantee that the correct trajectories and incorrect trajectories have the same total count for a given data point, which performs a simple 0-1 class balance. We train 1 epoch for filtered 76,792 samples with LoRA, where the rank $r$ and $\alpha$ are set to 64 and 128, respectively.
F.3 Evaluation
For evaluation, the temperature is set to $1.0$ and top_p is set to $0.9$. The max generation tokens are set to 8,192. Among the three presented decoding mechanisms, CoT (Wei et al., 2022) directly prompts models to generate a long reasoning chain and outputs the answer with “The answer is {answer}” for the convenience of answer extraction. Self-Consistency (Wang et al., 2023) generates $N=16$ samples for a given problem, and we select the one whose answer appears most times among the $N$ outputs. For PRM-guided Vote-Sum, we generate $N=12$ samples for datasets where SC has already achieved a high score (MedMCQA, Med-MMLU and BioASQ) and generate $N=16$ samples for other datasets. This is reasonable since not all generated samples are of high quality due to the sampling parameters top_p. We manually filter some low-probability samples to prevent them from disturbing the Vote-Sum computation. We use exact match (EM) to measure the performance. Specifically, we extract the contents following the last “The answer is” template to match the self-reflection thinking style, and perform appropriate post-process to derive a final prediction. For multiple-choice problems, we directly choose the first character of prediction phrases and measure whether the ground truth is equal to the prediction. For close-ended generation tasks, we remove quotes and turn the prediction and the ground truth into lowercase phrases. After that, we check whether the ground truth phrases exist in the prediction phrases.
F.4 Training Details of Distillation and RL
In this section, we elucidate the implementation details of distillation and RL.
Distillation
[…]
RL
We follow Guo et al. (2025) to use Group Relative Policy Optimization (GRPO; Shao et al. (2024)) to conduct RL training. We set the number of generations to 10 and the learning rate to $61e-$ 6. We adopt ZeRO-3 (Rajbhandari et al., 2020) to save memory and conduct full fine-tuning in one 8xA100 machine. The batch size is set to 4 per GPU. For the adopted prompt, we use the same prompt illustrated in DeepSeek-R1-zero, and use <think></think><answer></answer> to learn the slow-thinking output style. We use accuracy reward and format reward, and the reward setting is presented below in Table 6.
Table 6: Reward configurations for training RL-style medical models. Null means that there is no contents between <answer> and </answer>.
Conditions Accuracy Format
Correct 1 1
Incorrect -1 -1
Null 0 Null
Appendix G AI Assistant Usage
We mainly use AI assistants to polish our expressions.
Appendix H Case Study
Figure 8: Cases of MedQA
Figure 9: Cases of DDXPlus
Figure 10: Cases of SEER